The sale price of an item is $545.00. If bought on sale for 25% off what was the original price of the item?
Answer by Maths68(1474)

The sale price of an item is $545.00. If bought on sale for 25% off what was the original price of the item?
Let
Original price = x
25% of original price = x * 25% = x * 25/100 = x * 1/4 = x/4
Sale price = 545
Original price = Sale price + 25% of original price
x = 545 + x/4
Multiply by 4 both sides of the equation
4*x = 4*545 + 4*x/4
4x=2180+x
4x-x=2180
3x=2180
Divide by 3 both sides of above equation
3x/3=2180/3
x=726.66
Original price = x = 726.66
25% of original price = x * 25% = x * 25/100 = x * 1/4 = x/4 = 726.66/4 = 181.66
Sale price = 545